OpenStudy (anonymous):

Find the perimeter of an object which has a length of square root of 180 feet and a width of square root of 125 feet. Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. What type of object might this be? I've been working on this question for a while, would like help.

jhonyy9 (jhonyy9):

so first of all i think you need to know what type of object is this for you can calculi the perimeter you need to know how many sides with length and with has - how you have wrote above - i think not is possibly make the calculi

OpenStudy (anonymous):

You are given two dimensions, length and width, which suggests this is a rectangle. The perimeter of a rectangle is 2*width + 2*length

OpenStudy (anonymous):

Do you agree?

OpenStudy (anonymous):

2 * sqrt(180) + 2 * sqrt(125) = 2 * sqrt(5*36) + 2 * sqrt(5 * 25) = 2 * 6 * sqrt(5) + 2 * 5 * sqrt(5) = 12*sqrt(5) + 10*sqrt(5) = 22*sqrt(5)